OpenStudy (anonymous):

derivative of 4e^e^x

OpenStudy (anonymous):

4e^(x+e^x)

OpenStudy (abb0t):

Chain rule.

OpenStudy (abb0t):

\[4e^{e^x}\]

OpenStudy (anonymous):

4e^(e^x)e^x

OpenStudy (anonymous):

im confused now

OpenStudy (abb0t):

For this problem, you are required to use chain rule, which is: \[F'(x) = f'(g(x))g'(x)\] or \[\frac{ dy }{ dx } = \frac{ dy }{ du }\frac{ du }{ dx }\]

OpenStudy (abb0t):

1. factor out the constant (4) 2. take the derivative of e (which is also like f(x) 3. take the derivative of the inside function ( which is like g(x)) where f(x) = e and g(x) = e^x

OpenStudy (anonymous):

oh i see now, so basically it stays the same?

OpenStudy (abb0t):

Yes, but you have to also multiply it by the g(x) as you can see by the chain rule. So it kinda stays the sme but also looks slightly different.